OpenStudy (anonymous):

|x|+5=18 (1 point) 5 or - 5 13 or - 13 * (its this one right) 18 or - 18 23 or - 23

OpenStudy (anonymous):

Yes you are correct

OpenStudy (diyadiya):

13 or -13

OpenStudy (anonymous):

second yeah

OpenStudy (anonymous):

Thank you:)

OpenStudy (diyadiya):

x=18-5=13 -x=18-5=13 so x=-13

myininaya (myininaya):

|x|=x if x>0 |x|=-x if x<0 we need to solve equation for both cases x+5=18 or -x+5=18 x=13 -x=13 x=-13